Sol Summary – December 2nd

Crew 216 Sol Summary Report 02-DEC-19

Sol: 8

Summary Title: Communications and Filming Projects

Author: Marc Levesque

Mission Status: Active

Sol Activity Summary: Two EVAs were conducted in support of ongoing projects. In the morning Rich, Andrew, and Marc climbed to the summit of the North Ridge to confirm and map an accessible route and a location for a possible radio repeater site, both of which were accomplished. In the afternoon, Andrew, Michael, and Evgenia drove to the end of Quarry Road and then proceeded on foot for a short ways into Lithe Canyon for the purpose of capturing drone and documentary film footage.

Look Ahead Plan: Two EVAs are planned in support of Mike’s radioactive mineral project.

Anomalies in work: N/A.

Weather: Clear and cold.

Crew Physical Status: Healthy

EVAs: To North Ridge and Lithe Canyon.

Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, EVA Report, Astronomy Report, and Photos.
Support Requested: None

Research Report – December 1st

Crew 216 Astronomy Report 01-DEC-19
Crew Astronomer: Michael Ho

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: none
Objects to be Imaged this Evening: none
Images submitted with this report: nil
Problems Encountered: Waiting for 2 added observations that were requested from MDRS-14 since 25 and 27 Nov to be captured.

MUSK OBSERVATORY
Solar Features Observed: None.
Images submitted with this report: nil
Problems Encountered: Nil.

Research Report – December 1st

Crew 216 Research Report 01-DEC-19
Sol 7
Submitted by: Kennedy

1) Psych Scale (TAMU IRB2019-1042D) – 2nd stage evaluations were completed. The final will be done on the last day.

2) Medical Casts – Due to activities in the Hab no work was done on this project today.

3) In-Situ Fluorescent Mineral Prospecting – Planning reconnaissance EVA for Tues 03-Dec-19.

4) MDRS Mapping – Nothing further to report (awaiting return of map with desired annotations).

5) MDRS Communications – New ETA planned for north ridge reconnaissance. A new map with high-level detail was printed. As the last attempt to summit the north ridge for possible comms site was unsuccessful, we have decided to attempt the summit through a weakness in the south ridge.

EVA Report – December 1st

Crew 216 EVA Report 01-DEC-19
EVA # 09
Author: Whittle

Purpose of EVA: EVA footage for Roker film crew filming Crew 216

Start time: 1330
End time: 1515

Narrative: Crew left Hab in Curiosity and Spirit with Justin (Roker crew host) and crew. Arrived at Kissing Camel Ridge. Climbed high point on Kissing Camel Ridge at grid 12 S 0519008 4249540. Collected footage (camera and drone) for the Roker crew. Collected footage with rovers on way back to Hab.

Destination: Kissing Camel Ridge
Coordinates (use UTM NAD27 CONUS): 12 S 0519008 4249540.
Participants: Whittle, Kennedy, Ho, Justin + 5 film crew

Road(s) and routes per MDRS Map:
1) Hab to the point where Kissing Camel Ridge crossed Cow Dung Rd
2) Walked from Cow Dung Rd to the high point at 12 S 0519008 4249540
3) Walked from 12 S 0519008 4249540 back to rovers on Cow Dung Rd
4) Returned to Hab via Cow Dung Rd

Mode of travel: Curiosity, Spirit

Greenhab Report – December 1st

Crew 216 GreenHab Report 01-DEC-19
Crew GreenHab Officer: Michael Ho

Environmental control: Door opened for ventilation at 12.05 pm, closed at 7.50 pm.
Average temperature: 15C
Hours of supplemental light: Automatic system
Daily water usage for crops: 14L
Daily water usage for research and/or other purposes: nil
Water in Blue Tank – enough
Time(s) of watering for crops: 09.05 am; 4.15 pm.

SEASONAL PLANTING HISTORY:
(Established plants introduced since 03/11/2019 till 22/11/2019 please refer to Crew 215 GreenHab Report of 22/11/2019.)
Crew 216:
26/11/2019:
1 avocado seed in a small black pot
24/11/2019:
1 avocado seed in a small black pot
23/11/2019:
seeded Zucchini in 2 large black pots
seeded Chilli Pepper ‘Shishito’ in 2 2-gallon black pots
seeded Sweet Greenfield Pepper ‘California Wonder’ in 2 2-gallon black pots
seeded Swiss Chard ‘Flamingo Pink’ in 4 2-gallon black pots
transplanted three self-seeded tomato plants in large blue tub transplanted to 2-gallon pots

OTHER NOTES:
Adjusted the vents of the heater to blow horizontally instead of downwards. Removed temporary shades for strawberries as a result.
Will also monitor snow peas’ well-being as a result of an adjustment of the vents of the heater.
Harvest: None
Support/supplies requested:nil.

Commander Report – December 1st

Crew 216 Commander Report 01-DEC-2019
Sol: 7
Summary Title: Media visit
Author: Marc Levesque

Crew 216 became potential celebrities today when the Roker film crew spent the day filming for a science series that they hope to see aired next spring. Their filming of the MDRS facilities, crew interviews, and an EVA occupied us all day long.

The episode host, Justin, is a Ph.D. student at Columbia University who promotes STEM activities and studies, mostly to Gen Z students, but also to other groups. Mike gave an overview of the engineering aspects of our facilities, Michael provided a guided tour of the GreenHab and Observatory, and Andrew showed off the Science Dome.

Later Rich led an EVA that included Justin suited up as we do. Towards the end of the day, they did three individual interviews with Rich, Evgenia, and me, with each of us facing the camera and answering questions from Jacks, the director, who was standing with her back to us so as to not distract our looking at the camera. I felt a little awkward talking to the camera rather than a person, as I did the other day when Evgenia interviewed me but became comfortable after a while. Overall the film crew was very gracious, and we tried to provide them with the footage they wanted.

After they left, one crew member commented that their presence felt a bit intrusive to our sim, but I think all the crew realizes that the film should provide some great PR for the Mars Society and its MDRS and encourage more people to come here.

Operations Report – December 1st

Crew 216 Operations Report 01-DEC-19
SOL: 7
Name of person filing report: Mike Lawson

Non-nominal systems: NA

Generator: run
Hours run: 11.75
From what time last night: 1917
To what time this morning: 0656
List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before generator is run at night: 61 %

Diesel Reading – 37
Station Propane Reading – 55
Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 47
Water Meter (units): 1463701
Water (static tank) (gal): 292
Static to Loft Pump used – yes
Water in Green Hab (gal): 259
Water in Science Dome (gal): 0

Toilet tank emptied: no

Deimos rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used:
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: yes
Hours: 114.9
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 57.4%
Currently charging:

Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Curiosity rover used: yes
Hours: 119.9
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 56.0%
Currently charging: yes

Notes on rovers: Opportunity to be moved off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:

HabCar used and why, where?
CrewCar used and why, where?

General notes and comments: EVA to Kissing Camel Ridge. TV crew filmed.

Summary of internet: No serious issues.
Summary of suits and radios: No issues.
Summary of Hab operations: Propane restored to HAB kitchen. Water system pressure pump serviced.
Summary of GreenHab operations: routine watering.
Summary of Science Dome operations: TV crew filmed.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved.
Summary of any observatory issues: Sun observations performed. TV crew filmed.
Summary of health and safety issues: NA

Questions, concerns and requests to Mission Support: 2 bag NB trash, 1 bag expended dry cell batteries for recycling in the rear airlock. Expended 1 can of helmet screen cleaner, request refill.

Sol Summary – December 1st

Crew 216 Sol Summary Report 01-DEC-19
Sol: 7
Summary Title: Roker Film Crew
Author: Marc Levesque

Mission Status: Active

Sol Activity Summary: The crew spent the day hosting a film crew from Al Roker Entertainment. They were here to film one episode of a six-part series that will focus on different aspects of science. Upon arrival, they first did some filming outside the MDRS before joining us inside for some interviews and filming the Hab, GreenHab, Observatory, Science Dome, and RAM. The afternoon was spent filming an EVA with the episode host and then interviewing some individual crew members.

Look Ahead Plan: Two EVAs are planned to continue work on crew projects.

Anomalies in work: N/A.

Weather: Clear and cold.

Crew Physical Status: Healthy

EVA: To Kissing Camel Ridge for the Roker film crew.

Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, EVA Report, Astronomy Report, and Photos.

Support Requested: None

Journalist Report – December 1st

Crew 216 Journalist Report 01-DEC-19
Sol 6
Author: Evgenia Alexandrova

Solitude VS. Intimacy

We suppose that living on Mars will imply a lack of intimacy: you are closed up in a very limited space with strangers or close to those. Our commander Marc has experience in isolation at the American polar station in Antarctica. He states that “the confined setting tends to force people away from each other”.

Paradoxically, the lack of intimacy creates solitude. And solitude can drive people mad. To preserve astronauts from feeling lonely, we need to create enough psychological space for them. So far ARES 216 crew managed very well in establishing a comfortable and respectful environment for every crew member.

Today we had a TV crew visiting and filming. Strangely their presence felt like an intrusion. We’ve witnessed our intimacy being shaken. This is probably the best proof that we managed to see MDRS as our home.

Anonymous haiku of the day:

A stark red vista
Compelling, yet alien
Home is far away